A fast pitch softball player does a "windmill" pitch moving her hand through a vertical circular arc to pitch a ball at 70mph. The 0.20 kg ball is 50 cm from the pivot point at her shoulder.

Part A Just before the ball leaves her hand, what is its centripetal acceleration? Express your answer in meters per second squared.

Part B At the lowest point of the circle, the ball has reached its maximum speed. What is the magnitude of the force her hand exerts on the ball at this point? Express your answer in newtons.

Part C At the lowest point of the circle, the ball has reached its maximum speed. What is the direction of the force her hand exerts on the ball at this point? (upward, downward, left, or right)
Answer #1

Convert speed, v = 70 mph in m/s.

v = 70 mph = (70 x 1609) / (60 x 60) = 31.3 m/s

Part A -

Radius, R = length of the shoulder = 50 cm = 0.50 m

Centripetal acceleration, a = v^2 / R = 31.3^2 / 0.50 = 1958.0 m/s^2

Part B -

Mass of the ball, m = 0.20 kg

Magnitude of force, the player's hand exerts on the ball, F = m*a = 0.20 * 1958 = 391.6 N (Answer)

Part C -

The force exerted by player's hand on the ball is in the upward direction.